OpenStudy (anonymous):

Triangle ABC is similar to triangle XYZ.The length AB is 10.The length BC is 16.Find the length XY if the length YZ is 12.

OpenStudy (anonymous):

10/x=16/12

OpenStudy (anonymous):

Then cross multiply to find the value of x, which equals the length of XY.

OpenStudy (anonymous):

can you show me please

OpenStudy (anonymous):

10/x=16/12, so 16x=12(10). 16x=120 x=120/16, so x=7.5